# Output Range Analysis for Deep Neural Networks based on Simulated Annealing Processes

### v1

@@ Abstract
This paper tackles the challenging problem of output range estimation for Deep Neural Networks (DNNs), introducing a novel algorithm based on Simulated Annealing (SA). Our approach addresses the lack of local geometric information and high non-linearity in DNNs, making it versatile across various architectures, especially Residual Neural Networks (ResNets). We present a straightforward, implementation-friendly algorithm that avoids restrictive assumptions about network architecture. Through theoretical analysis and experimental evaluations, including tests on the Ackley function, we demonstrate our algorithm’s effectiveness in navigating complex, non-convex surfaces and accurately estimating DNN output ranges. Futhermore, the Python codes of this experimental evaluation that support our results are available in our GitHub repository***https://github.com/Nicerova7/output-range-analysis-for-deep-neural-networks-with-simulated-annealing.
Keywords Deep Neural Networks, Simulated Annealing, Output Range Estimation, Residual Neural Networks, Global Optimization, Cyber-Physical Systems, Non-Convex Optimization
@@ 1 Introduction
Unquestionably, in recent decades, Deep Neural Networks (DNNs) have been by far the most widely used tools to perform complex machine learning tasks. More recently, DNNs have been used in cyber-physical systems critical to public security and integrity; such as autonomous vehicle driving and air traffic systems. Therefore, it is of pressing interest to implement security verification systems for DNNs. One of the objectives in this line of interest is the verification of the maximum and minimum values assumed by a DNN, an objective commonly known as the range estimation problem, see Dutta et al., (2018); Wang et al., (2018). However, the relationships established between the inputs and outputs of a DNN are highly non-linear and complex, difficult to understand with existing tools today. Due to this inability, DNNs are commonly referred to as black boxes. This nature of DNN makes the range estimation problem particularly challenging, because there is no geometric information about the response surface generated by a DNN. For example, if local geometric information about the generated surface were obtained, such as the gradient vector and the Hessian matrix at each point, the problem could be addressed with conventional nonlinear programming techniques. However, in a DNN it is only possible to obtain point information about the estimated response, without any local knowledge around that point. These facts, added to the high non-linearity of a DNN, make the range estimation problem non-trivial. Recently, various methodologies have been proposed for the analysis of the output range of DNNs, see for example Dutta et al., (2018); Wang et al., (2018); Tran et al., (2020); Huang et al., (2020); Liu et al., (2021). However, these methodologies make restrictive assumptions about the network architecture which limits the use and application of these methodologies. In this paper, based on global optimization techniques, in particular Simulated Annealing, we present a simple algorithm, which contemplates the lack of local geometric information of the surfaces generated by DNNs, easy to implement and use, and that solves the range estimation problem for a wide spectrum of neural networks in general.
@@
Outline. In Section 2 we make a mathematical description of a very particular family of deep neural networks called Residual Networks. Although our methodology works for any particular family, we focus on these neural networks given their relevance in applications and for illustrative purposes of the use of our algorithm. In Section 3 we specifically define the output range analysis problem for neural networks, mentioning its characteristics, limitations and challenges. In Section 4 we develop an exhaustive treatment of Simulated Annealing for limited domains, adapting its properties and guaranteeing its results for our purposes. In Section 5, from the above, we derive an algorithm to solve the output range analysis problem. Finally, in Section 6 we present a part of the experimental evaluations that we have executed to numerically guarantee our results and the performance of our proposed algorithm.
@@ 2 Residual Neural Networks
As already mentioned before, although our algorithm works for a wide spectrum of neural networks in general, we focus on Residual Neural Networks given their relevance in applications and for illustrative purposes of the use of our algorithm. The Residual Networks (or ResNets) were introduced in He et al., (2016), for which in this chapter we show a mathematical description. Let $x\in\mathbb{R}^{d}$ denote the inputs and $y\in\mathbb{R}$ be the output in a supervised learning problem. In this learning context, the objective is the approximation of a function $f:x\in\mathbb{R}^{d}\mapsto y\in\mathbb{R}$ using a ResNet, which we denote as $\mathcal{F}$. The basic unitary components of an ResNet are called neurons, which are organized in a consecutive number of layers that are linked through nonlinear functions. To obtain a mathematical description of the operations carried out within a ResNet, we define some notations. We consider a network with a source layer, a output layer and $L$ hidden layers, where the $l$-th layer contains $H_{l}$ neurons ($l$-th layer width), for $l=0,\dots,L+1$. Note that $H_{0}=d$ and $H_{L+1}=1$, corresponding to the feature inputs and target outputs respectively. We denote the output vector of the $l$-th layer by $x^{(l)}\in\mathbb{R}^{H_{l}}$ which corresponds to the input of the next layer. We set $x^{(0)}=x\in\mathbb{R}^{d}$ and $x^{(L+1)}=y\in\mathbb{R}$. For $l=1,\dots,L$, the $i$-th neuron performs an affine transformation on that layers input $x^{(l-1)}$ followed by a non-linear transformation
@@
Although the activation function can, in principle, be chosen arbitrarily, there are three functions that have particularly proven to be useful in various applications; ReLU, Sigmoid and Tanh, see LeCun et al., (2015); Goodfellow et al., (2016). The parameters of the network is all the weights and biases $\{W^{(l)},b^{(l)}\}_{l=1}^{L+1}\in\mathbb{R}^{N_{p}}$, where $N_{p}=\sum\limits_{l=1}^{L+1}(H_{l-1}+1)H_{l}$ is the total number of parameters.
@@ 3 Output Range Analysis Problem
    INPUTS: : Neural network, for example a ResNet $\mathcal{F}$, and input constraints $Ax\leq b$ which generates a feasible set $E$, i.e., $E=\{x\in\mathbb{R}^{d}:Ax\leq b\}$.
@@
Due to the applications, we focus just on hypercubes of $\mathbb{R}^{d}$ as feasible set, that is; sets of the type $E=[l_{1},u_{1}]\times\cdots\times[l_{d},u_{d}]\subset\mathbb{R}^{d}$, where $[l_{j},u_{j}]$ are intervals in $\mathbb{R}$.
As mentioned above, this problem is particularly challenging given that there is generally almost total ignorance about the surfaces generated by deep neural networks. That is, there is no information on the functional form of the mapping $\mathcal{F}$. Therefore, there is no local geometric information such as the local slopes and curvatures of these surfaces, making the use of traditional nonlinear programming methodologies impossible. Furthermore, the surfaces generated by this type of neural networks are strictly non-convex, with many crazy minima, therefore it requires thinking about global optimization methods, methods such as Simulated Annealing.
@@ 4 Simulated Annealing with Boundary Conditions
Our range estimation problem can be summarized in an optimization problem of the form
@@
where $\mathcal{F}$ is a ResNet and $E$ is a hypercubes of $\mathbb{R}^{d}$, which we denote by the pair $(\mathcal{F},E)$. The main problem now is to find a point $x_{\min}\in E$ such tha $\mathcal{F}(x_{\min})$ is global minimal on $E$. The $\max_{x\in E}\mathcal{F}$ is equivalent to $\min_{x\in E}-\mathcal{F}$, which is why we can, without loss of generality, only talk about “Minimization” throughout this paper. We denote by $M_{\mathcal{F}}=\{x_{\min}\in E:\mathcal{F}(x_{\min})\leq\mathcal{F}(x),\,\textrm{for all}\,\,x\in E\}$ the minimum set of $\mathcal{F}$. From Equation (3), we know that $\mathcal{F}$ is a continuous function, and added to the fact that $E$ is a compact set, we know that $M_{\mathcal{F}}$ is a non-empty set. For simplicity in notation we set $\mathcal{F}_{\min}=\mathcal{F}(x_{\min})$.
Definition 4 (Reflection).
@@
[…]
@@
where $B\subset\mathfrak{B}=\mathfrak{B}(E)$, and
@@
$q_{T}(x,y)=e^{\frac{1}{T}\min\{0,\mathcal{F}(x)-\mathcal{F}(y)\}}.$ (8)
Here $T$ is called temperature parameter.
@@
$P_{i}(x,B)=\begin{cases}\int\limits_{B}Q_{i}(x,\mathrm{d}y)\,q_{T_{i}}(x,y)\,\mathrm{d}y&\text{for }x\notin B,\\ \int\limits_{B}Q_{i}(x,\mathrm{d}y)\,q_{T_{i}}(x,y)\,\mathrm{d}y+\Big{(}1-\int\limits_{E}Q_{i}(x,\mathrm{d}y)\,q_{T_{i}}(x,y)\,\mathrm{d}y\Big{)}&\text{for }x\in B,\end{cases}$ (9)
@@
$\int\limits_{B}{Q}(x,\mathrm{d}y)=\int\limits_{B}{Q}(y,\mathrm{d}x),$ (12)
@@ 5 Algorithm Derivation
From Theorem 8 and its Corollary 9, a simple algorithm can easily be derived to solve the range estimation problem $(\mathcal{F},E)$. The main idea is to generate a Markov process $(X_{i})_{i\in\mathbb{N}}$ taking values on $E$, as described in Definition 7, for which its initial state is selected using an arbitrary measure $\mu_{0}$. The following states of the Markov process are selected according to two stages; In the first stage, a new state generated by $Q$ is proposed, and subsequently in the second stage, it is decided to accept or reject that new state according to the probability of acceptance $q$. If the proposed state is outside the domain $E$ we use reflection $\mathcal{R}$ to exchange it for another equally probable state that is inside the domain $E$. It is important to note that the generation of new proposed states using $Q$ is done through $\rho$. Both stages, proposing and accepting new states, are executed repeatedly $N$ times for each temperature level $T$. Theorem 8 guarantees that if we execute this two-stage local recursive process the Markov process $(X_{i})_{i\in\mathbb{N}}$ will converge towards the minimal state corresponding to the value $\mathcal{F}_{\min}$, this occurs for each set temperature level. Finally, Corollary 9 guarantees that if we execute the stages described above for a sequence of temperatures $\{T_{i}\}_{i\in\mathbb{N}}$ that slowly decreases towards zero, then the Markov process $(X_{i})_{i\in\mathbb{N}}$ will converge to the minimal state corresponding to the global minimum that solves our estimation problem. We summarize the entire procedure described above in pseudocode format in Algorithm 1. In addition, we provide a Python code that is available in https://github.com/Nicerova7/output-range-analysis-for-deep-neural-networks-with-simulated-annealing.
Algorithm 1 Minimum Output Search for a Deep Neural Network
Input: The minimization problem $(\mathcal{F},E)$
Initialize $i\leftarrow 0$, $x\sim\mu_{0}$, $T_{0}\leftarrow T_{\max}$ and $\mathcal{F_{\min}}\leftarrow\mathcal{F}(x)$
while $T_{i}>T_{\min}$ do
     for $k=1$ to $N$ do
         Generate $y\sim\rho(x,y)$
         Reflect $y\leftarrow\mathcal{R}(y)$
         Set $\Delta\mathcal{F}\leftarrow\mathcal{F}(y)-\mathcal{F}(x)$
         if $\Delta\mathcal{F}<0$ then
              $q(x,y)=1.0$
         else
              $q(x,y)=\min\Big{\{}1,e^{-\frac{\Delta\mathcal{F}}{T_{i}}}\Big{\}}$
         end if
         Sample $U\sim\textrm{Unif}\,(0,1)$
         if $U\leq q(x,y)$ then
              Update $x\leftarrow y$
         end if
         if $\mathcal{F}(x)<\mathcal{F}_{\min}$ then
              $\mathcal{F}_{\min}\leftarrow\mathcal{F}(x)$
         end if
     end for
     Update $T_{i}=T_{i-1}\,\delta^{i}$
     Update $i\leftarrow i+1$
end while
Output: $\mathcal{F}_{\min}$
6 Experimental Evaluation

### v3

# A simple algorithm for output range analysis for deep neural networks

@@ Abstract
This paper presents a simple algorithm for the output range estimation problem in Deep Neural Networks (DNNs) by integrating a Simulated Annealing (SA) algorithm tailored to operate within constrained domains and ensure convergence towards global optima. The method effectively addresses the challenges posed by the lack of local geometric information and the high non-linearity inherent to DNNs, making it applicable to a wide variety of architectures, with a special focus on Residual Networks (ResNets) due to their practical importance. Unlike existing methods, our algorithm imposes minimal assumptions on the internal architecture of neural networks, thereby extending its usability to complex models. Theoretical analysis guarantees convergence, while extensive empirical evaluations—including optimization tests involving functions with multiple local minima—demonstrate the robustness of our algorithm in navigating non-convex response surfaces. The experimental results highlight the algorithm’s efficiency in accurately estimating DNN output ranges, even in scenarios characterized by high non-linearity and complex constraints. For reproducibility, Python codes and datasets used in the experiments are publicly available through our GitHub repository ***https://github.com/Nicerova7/output-range-analysis-for-deep-neural-networks-with-simulated-annealing.
Keywords Output Range Analysis, Deep Neural Networks, Simulated Annealing, Global Optimization, Non-Convex Optimization, Residual Networks.
@@ 1 Introduction
Unquestionably, in recent decades, Deep Neural Networks (DNNs) have been by far the most widely used tools to perform complex machine learning tasks. More recently, DNNs have been used in cyber-physical systems critical to public security and integrity; such as autonomous vehicle driving and air traffic systems. Therefore, it is of pressing interest to implement security verification systems for DNNs. One of the objectives in this line of interest is the verification of the maximum and minimum values assumed by a DNN, an objective commonly known as the range estimation problem, see Dutta et al., (2018); Wang et al., (2018). This interest in estimating the range assumed by a DNN responds to the objective of diagnosing and validating the previously executed training. However, the relationships established between the inputs and outputs of a DNN are highly non-linear and complex, making it difficult to understand with existing tools today. Due to this inability, DNNs are commonly referred to as black boxes. This nature of DNN makes the range estimation problem particularly challenging because there is no geometric information about the response surface generated by a DNN. For example, if local geometric information about the generated surface by a DNN was obtained, such as the gradient vector and the Hessian matrix at each point, the problem could be addressed with conventional nonlinear programming techniques. However, in a DNN it is only possible to obtain point information about the estimated response, without any local knowledge around that point. These facts, added to the high nonlinearity of a DNN, make the range estimation problem nontrivial. In particular, this problem is even more challenging when dealing with a DNN, in contrast to other types of neural network, since its multiple layers increase the complexity of the problem. In this paper, we present a simple algorithm based on global optimization techniques, initially motivated by the classical Simulated Annealing (SA) kirkpatrick1983optimization, to solve the range estimation problem for a wide spectrum of neural networks, and in particular DNNs. In contrast to classical SA, our algorithm considers the existence of a restricted domain for searching for optimal points, which eventually corresponds to the domain of the training data of the analyzed neural network. Our algorithm does not make use of any information about the internal architecture of the analyzed DNN, which makes it applicable to a much larger spectrum than other proposals available in the literature, see for example Dutta et al., (2018); Wang et al., (2018); katz2017reluplex. Our algorithm considers restricted search spaces, which underlie the nature of the output range estimation problem of neural networks. Furthermore, we present results, both theoretical and empirical, that guarantee the convergence of our algorithm towards the optimal points, which leads to a good estimation of the output range.
@@
Outline. In the section “Residual Neural Networks" we make a mathematical description of a very particular family of deep neural networks called Residual Networks. Although our methodology works for any particular family, we focus on these neural networks given their relevance in applications and for illustrative purposes of the use of our algorithm. In the section “Output Range Analysis Problem", we specifically define the output range analysis problem for neural networks, mentioning its characteristics, limitations, and challenges. In the section titled “Simulated Annealing with Boundary Conditions" we develop an exhaustive treatment of Simulated Annealing for limited domains, adapting its properties and guaranteeing its results for our purposes. In the section titled “Algorithm Derivation", from the above, we derive an algorithm to solve the output range analysis problem. Finally, in the section “Experimental Evaluation" we present a part of the experimental evaluations that we have executed to numerically guarantee our results and the performance of our proposed algorithm.
2 Related Work
Recently, various methodologies have been proposed for the analysis of the output range of DNNs. Dutta et al., (2018) propose a novel approach to estimate the output ranges of these models given specific input sets. Their novel algorithm utilizes local search techniques and linear programming to efficiently compute the maximum and minimum values in a DNN over the input set. The algorithm repeatedly applies local gradient descent to identify and eliminate local minima of the neural network’s output function. Once local optima are identified, the final global optimum is validated through a mixed integer programming model. The authors consider the algorithm’s computational efficiency while still maintaining the accuracy of the output range estimation. However, common issues like scalability and the potential complexity of a DNN are not fully addressed. Furthermore, the methodology makes specific assumptions about the neural network, which means that the methodology can only be applied to a rather restricted class of neural networks.
On the other hand, autonomous vehicles, collision avoidance systems, and others are part of a real-world security-critical domain, making it important to check the security properties of DNNs. These security properties guarantee that a DNN should have to ensure it operates safely and reliably against potential threats. There are frameworks for verifying the security properties of DNN that use interval arithmetic Wang et al., (2018); katz2017reluplex; carlini2018provably. Wang et al., (2018) shows us a framework called ReluVal, which employs interval arithmetic to represent the ranges of inputs and outputs. For each input and operation in a DNN, the interval calculation is performed to explore the output range of the network. By analyzing these intervals, the framework can detect inputs that lead to incorrect classifications or outputs, thereby exposing potential weaknesses in the network. The experiments demonstrate the efficiency of ReluVal in verifying security properties, outperforming other frameworks. Nevertheless, it may still struggle with scalability when applied to extensive networks or complex architectures. Furthermore, the framework is focused on specific types of adversarial attacks, leaving uncertainty regarding its performance on different architectures.
Other kinds of verification of DNNs are required: safety and robustness. Tran et al., (2020) presents the Neural Network Verification (NNV) software tool. The tool applies a collection of reachability algorithms and the use of a variety of set representations, such as polyhedra, star sets, zonotopes, and abstract-domain representations. Also, NNV can handle different types of neural networks and system models, allowing it to support both exact and over-approximate analysis. However, representations as polyhedra are limited by scalability and may not work effectively, as they involve costly operations, leading to a conservative and less practical reachability analysis in larger systems. Moreover, all experiments focus on specific exercises, which limits the exploration of new and complex DNN architectures. Then, Liu et al., (2021) provide a comprehensive overview of methods developed to formally verify the safety and robustness of DNNs. The authors discuss different algorithms that incorporate reachability analysis, optimization, and search techniques to verify the appropriate configuration of the DNNs for the input-output properties across their entire input space. These algorithms are classified into different categories, such as layer-by-layer, reachability analysis, optimization-based, and combined search and verification methods. However, after exhaustive verification of these algorithms, the work concludes with some algorithms that present, like the other works, problems with the network scale. Finally, Huang et al., (2020) analyze the output range of a neural network using a convex polygonal relaxation (over-approximation) of the activation functions to manage nonlinearity, allowing the problem to be formulated as a mixed-integer linear program (MILP). However, a key limitation is the addition of more integer variables in the MILP, which increases computational complexity, particularly in DNNs or those with more complex architectures.
In general, these proposals described above make restrictive assumptions about the network architecture which limits the use and application of these methodologies. In contrast, our proposal does not make any assumptions about the internal architecture of the evaluated neural network, which makes our algorithm widely applicable in output rank analysis for neural networks, and given the greater complexity, even more so when it comes to DNNs. Given these features of our algorithm, added to this are the theoretical and empirical guarantees that we present, and its simple and easy implementation, we believe that it is a relevant and practical contribution to the objectives in this application area.
@@ 3 Residual Neural Networks
As mentioned above, since we make no assumptions about the internal architecture of the analyzed neural network, our algorithm works for a wide spectrum of neural networks. However, without loss of generality, we focus on Residual Neural Networks given their relevance in applications and for illustrative purposes of the use of our algorithm. The Residual Networks (or ResNets) is a kind of DNN that was introduced in He et al., (2016), for which in this chapter we show a mathematical description. Let $x\in\mathbb{R}^{d}$ denote the inputs and $y\in\mathbb{R}$ be the output in a supervised learning problem. In this learning context, the objective is the approximation of a function $f:x\in\mathbb{R}^{d}\mapsto y\in\mathbb{R}$ using a ResNet, which we denote as $\mathcal{F}$. The basic unitary components of a ResNet are called residual blocks, which are organized in a consecutive number of layers that are linked through nonlinear functions. To obtain a mathematical description of the operations carried out within a ResNet, we define some notations. We consider a network with a source layer, a output layer and $L$ hidden layers, where the $l$-th layer contains $H_{l}$ neurons ($l$-th layer width), for $l=0,\dots,L+1$. Note that $H_{0}=d$ and $H_{L+1}=1$, corresponding to the feature inputs and target outputs respectively. We denote the output vector of the $l$-th layer by $x^{(l)}\in\mathbb{R}^{H_{l}}$ which corresponds to the input of the next layer. We set $x^{(0)}=x\in\mathbb{R}^{d}$ and $x^{(L+1)}=y\in\mathbb{R}$. For $l=1,\dots,L$, the $i$-th neuron performs an affine transformation on that layers input $x^{(l-1)}$ followed by a non-linear transformation
@@
Although the activation function can, in principle, be chosen arbitrarily, there are three functions that have particularly proven to be useful in various applications; ReLU, Sigmoid, and Tanh, see LeCun et al., (2015); Goodfellow et al., (2016). The parameters of the network is all the weights and biases $\{W^{(l)},b^{(l)}\}_{l=1}^{L+1}\in\mathbb{R}^{N_{p}}$, where $N_{p}=\sum\limits_{l=1}^{L+1}(H_{l-1}+1)H_{l}$ is the total number of parameters.
@@ 4 Output Range Analysis Problem
    INPUTS: : A trained neural network (for example a ResNet $\mathcal{F}$) and input constraints $Ax\leq b$ which generates a feasible set $E$, i.e., $E=\{x\in\mathbb{R}^{d}:Ax\leq b\}$.
@@
Due to the applications, we focus just on hypercubes of $\mathbb{R}^{d}$ as feasible set, that is; sets of the type $E=[l_{1},u_{1}]\times\cdots\times[l_{d},u_{d}]\subset\mathbb{R}^{d}$, where $[l_{j},u_{j}]$ are intervals in $\mathbb{R}$. Note that $E$ is the restricted search domain in which we want to diagnose and validate the analyzed neural network.
As mentioned above, this problem is particularly challenging given that there is generally almost total ignorance about the surfaces generated by a DNN. That is, there is no information on the functional form of the mapping $\mathcal{F}$. Therefore, there is no local geometric information such as the local slopes and curvatures of these surfaces, making the use of traditional nonlinear programming methodologies impossible. Furthermore, the surfaces generated by this type of neural network are strictly non-convex, with many minima. Therefore, it requires thinking about global optimization methods, methods such as SA for which it is necessary to incorporate the restricted search domain $E$.
@@ 5 Simulated Annealing with Boundary Conditions
Our objective in this session is to incorporate the domain $E$, coming from the range estimation problem, into the classical SA, and at the same time establish theoretical guarantees of the optimization process of the neural network $\mathcal{F}$. To this end, our range estimation problem can be summarized in an optimization problem of the form
@@
where $\mathcal{F}$ is a ResNet and $E$ is a hypercubes of $\mathbb{R}^{d}$, which we denote by the pair $(\mathcal{F},E)$. The main problem now is to find a point $x_{\min}\in E$ such tha $\mathcal{F}(x_{\min})$ is global minimal on $E$. The $\max_{x\in E}\mathcal{F}$ is equivalent to $\min_{x\in E}-\mathcal{F}$, which is why we can, without loss of generality, only talk about “Minimization” throughout this paper. We denote by $M_{\mathcal{F}}=\{x_{\min}\in E:\mathcal{F}(x_{\min})\leq\mathcal{F}(x),\,\textrm{for all}\,\,x\in E\}$ the minimum set of $\mathcal{F}$. From Equation (3), we know that $\mathcal{F}$ is a continuous function, and added to the fact that $E$ is a compact set, we know that $M_{\mathcal{F}}$ is a non-empty set. For simplicity in notation we set $\mathcal{F}_{\min}=\mathcal{F}(x_{\min})$. Next, we will incorporate the search domain by establishing reflective band conditions on $E$.
Definition 4 (Cyclic Reflection).
@@
$\mathcal{R}(y_{j})=\begin{cases}l_{j}+\big{[}(y_{j}-l_{j})\mod(u_{j}-l_{j})\big{]},&\text{if }(y_{j}-l_{j})\mod\big{(}2(u_{j}-l_{j})\big{)}\leq u_{j}-l_{j},\\[10.0pt] u_{j}-\big{[}(y_{j}-l_{j})\mod(u_{j}-l_{j})\big{]},&\text{if }(y_{j}-l_{j})\mod\big{(}2(u_{j}-l_{j})\big{)}>u_{j}-l_{j}.\end{cases}$ (5)
@@
where $B\subset\mathfrak{B}$, and $\mathcal{B}(E)$ denotes the Borel $\sigma$-algebra on the state space $E$.
@@
$q_{T}(x,y)=e^{\frac{1}{T}\min\{0,\mathcal{F}(x)-\mathcal{F}(y)\}},$ (8)
where $T$ is called temperature parameter.
@@
$P_{i}(x,B)=\begin{cases}\int\limits_{B}q_{T_{i}}(x,y)\,Q_{i}(x,y)\,\mathrm{d}y&\text{for }x\notin B,\\ \int\limits_{B}q_{T_{i}}(x,y)\,Q_{i}(x,y)\,\mathrm{d}y+\Big{(}1-\int\limits_{E}q_{T_{i}}(x,y)\,Q_{i}(x,y)\,\mathrm{d}y\Big{)}&\text{for }x\in B,\end{cases}$ (9)
@@
This definition aligns with the framework established in Definition 2.4 of Haario and Saksman Haario and Saksman, (1991), which introduces the simulated annealing process in a general state space. Their work provides a rigorous foundation for the transition kernels and the Markov process in this context.
@@
$\int\limits_{B}{Q}(x,y)\,\mathrm{d}y=\int\limits_{B}{Q}(y,x)\,\mathrm{d}x,$ (12)
@@
As will be seen later, Corollary 9 is the mechanism that establishes the theoretical guarantees that our algorithm converges to the global minimum point.
@@ 6 Algorithm Derivation
From Theorem 8 and its Corollary 9, a simple algorithm can easily be derived to solve the range estimation problem $(\mathcal{F},E)$. The main idea is to generate a Markov process $(X_{i})_{i\in\mathbb{N}}$ taking values on $E$, as described in Definition 7, for which its initial state is selected using an arbitrary measure $\mu_{0}$. The following states of the Markov process are selected according to two stages; In the first stage, a new state generated by $Q$ is proposed, and subsequently in the second stage, it is decided to accept or reject that new state according to the probability of acceptance $q$. If the proposed state is outside the domain $E$ we use reflection $\mathcal{R}$ to exchange it for another equally probable state that is inside the domain $E$. It is important to note that the generation of new proposed states using $Q$ is done through $\rho$. Both stages, proposing and accepting new states, are executed repeatedly $N$ times for each temperature level $T$. Theorem 8 guarantees that if we execute this two-stage local recursive process the Markov process $(X_{i})_{i\in\mathbb{N}}$ will converge towards the minimal state corresponding to the value $\mathcal{F}_{\min}$, this occurs for each set temperature level. Finally, Corollary 9 guarantees that if we execute the stages described above for a sequence of temperatures $\{T_{i}\}_{i\in\mathbb{N}}$ that slowly decreases towards zero, then the Markov process $(X_{i})_{i\in\mathbb{N}}$ will converge to the minimal state corresponding to the global minimum that solves our estimation problem. We summarize the entire procedure described above in pseudo-code format in Algorithm 1. In addition, we provide a Python code that is available in https://github.com/Nicerova7/output-range-analysis-for-deep-neural-networks-with-simulated-annealing.
Algorithm 1 Minimum Output Search for a Neural Network
1:Input: The minimization problem $(\mathcal{F},E)$
2:Initialize $i\leftarrow 0$, $x\sim\mu_{0}$, $T_{0}\leftarrow T_{\max}$ and $\mathcal{F_{\min}}\leftarrow\mathcal{F}(x)$
3:Output: $\mathcal{F}_{\min}$
4:while $T_{i}>T_{\min}$ do
5:     for $k=1$ to $N$ do
6:         Generate $y\sim\rho(x,y)$
7:         Reflect $y\leftarrow\mathcal{R}(y)$
8:         Set $\Delta\mathcal{F}\leftarrow\mathcal{F}(y)-\mathcal{F}(x)$
9:         if $\Delta\mathcal{F}<0$ then
10:              $q(x,y)=1.0$
11:         else
12:              $q(x,y)=\min\Big{\{}1,e^{-\frac{\Delta\mathcal{F}}{T_{i}}}\Big{\}}$
13:         end if
14:         Sample $U\sim\textrm{Unif}\,(0,1)$
15:         if $U\leq q(x,y)$ then
16:              Update $x\leftarrow y$
17:         end if
18:         if $\mathcal{F}(x)<\mathcal{F}_{\min}$ then
19:              $\mathcal{F}_{\min}\leftarrow\mathcal{F}(x)$
20:         end if
21:     end for
22:     Update $T_{i}=T_{i-1}\,\delta^{i}$
23:     Update $i\leftarrow i+1$
24:end while
25:Output: $\mathcal{F}_{\min}$
7 Experimental Evaluations
Based on the above constructions, our algorithm can be used to estimate the maxima and minima assumed by neural networks that were trained using an available database. It is important to mention that the database provided for training the neural network does not necessarily contain the maxima and minima assumed by this neural network. These maxima and minima correspond to the optimal points of the continuous response surface generated by the neural network that arises from the adjustment of the discrete observations contained in the training database. The response surface generated by a neural network can be viewed as a smoothing of the data. But, as mentioned above, we do not know the explicit form of the function graph that corresponds to the response surface generated by the trained neural network. Therefore, our strategy to empirically evaluate our algorithm consists in proposing some explicit functional forms for which we know their optimal points, in this case their minimum points. We will use these explicit functions, for which we have all possible geometric information, including their minimum points, to generate a noisy discrete sample through uniform sampling in its domain. These discretized samples are used as training data for a ResNet. The main idea of our experimental evaluation strategy is that by applying this algorithm to the ResNet, trained with the discretized data, we can approximate the true global minima of the functions that were used to generate these discretized samples.
Following the empirical evaluation scheme proposed above, and to illustrate the use and operation of our algorithm, we will use a couple of commonly used functions for testing optimization algorithms. This functions are: Ackley Function, Drop-Wave Function and a function with multiple global minima.
7.1 The Ackley Function